OpenStudy (anonymous):

Use logarithmic differentiation to calculate the derivative... e^2x multiplied by sqrt(1+x^2) divided by tan(x)^2

OpenStudy (anonymous):

\[2x+\log(\sqrt{1+x^2}))-2logtanx\]\[\frac{dy}{dx}=2+\frac{1}{\sqrt{1+x^2}}(2x)-2(\frac{1}{tanx})(\sec^2x)\]

OpenStudy (anonymous):

That's the answer i got but according to my teacher the \[2+ x \div 1+x ^{2}- 4\csc2x\]

OpenStudy (anonymous):

and thats multiplied by the original equation

OpenStudy (anonymous):

i think what Johnny has worked out is dy/d x * 1/y (where y = f(x)

OpenStudy (anonymous):

i answered your question in your last post lilon

OpenStudy (anonymous):

i've had another look at this lilon and the answer according to your teacher is right = as you say multiplied by the original equation dy/dx = e^2x[2 +2x^2 + x - 4(1 + x^2)csc2x] / (1+x^2)^1/2*tan(x)^2